40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL  Not logged in ELOG logo
Entry  Fri Feb 4 01:03:47 2011, Frank, DailyProgress, BEAT, cable delay setup cable-delay.png
    Reply  Fri Feb 4 13:59:48 2011, Frank, DailyProgress, BEAT, trend of the setup Untitled.png
Message ID: 470     Entry time: Fri Feb 4 01:03:47 2011     Reply to this: 471
Author: Frank 
Type: DailyProgress 
Category: BEAT 
Subject: cable delay setup 

changed the frequency noise readout to the cable-delay version to see how it works.
As the loss of the cable is very large and the signal from the photodetector not very strong i had to add some amplifiers for signal conditioning (see figure below).
Data is acquired with channel C3:PSL-FSS_FREQCOUNT in replacement for the frequency counter. Started taking data around 11/2/4 4:16:30 UTC.
Will do final calibration tomorrow. First test gave signal from mixer changes from peak-to-peak for about 600kHz (+/-10%) in frequency change.

cable-delay.png

ELOG V3.1.3-